OpenStudy (anonymous):

Solve for x. 3x − 1 = -10

OpenStudy (anonymous):

Simplifying 3x + -1 = -10 Reorder the terms: -1 + 3x = -10 Solving -1 + 3x = -10 Move all terms containing x to the left, all other terms to the right. Add '1' to each side of the equation. -1 + 1 + 3x = -10 + 1 Combine like terms: -1 + 1 = 0 0 + 3x = -10 + 1 3x = -10 + 1 Combine like terms: -10 + 1 = -9 3x = -9 Divide each side by '3'. x = -3 Simplifying x = -3

OpenStudy (anonymous):